16) A rollercoaster that is 20. m high starts going down the track with an initial velocity of 8.0 m/s and then approaches a loop with a radius of 3.0 m. What is the rollercoasters final velocity forces (friction, air resistance) are negligible and that the loop is a perfect circle. Vo- 8.0 m/s
Answer #1

Using energy conservation

mgh + 0.5 m u^2 = 0.5 m v^2 + mg ( 2R)

9.8* 20 + 0.5* 8^2 = 0.5* v^2 + 9.8* 2* 3

v = 18.396 m/s

=====

Comment in case any doubt please rate my answer... .